OpenStudy (anonymous):

A circle whose center is at (6,8) passes through the origin, Which of the following points is NOT on the circle?

hartnn (hartnn):

can u find the equation of circle ?

OpenStudy (anonymous):

well

OpenStudy (anonymous):

I use slope formula

OpenStudy (anonymous):

8-0/6-0=8/6

OpenStudy (anonymous):

what's the equation of a circle?

OpenStudy (anonymous):

I forgot...

hartnn (hartnn):

standard form : (x-h)^2 + (y-k)^2 = r^2 centre = (h,k) radius =r

OpenStudy (anonymous):

ahh

OpenStudy (anonymous):

so I could not just use distance formula?

hartnn (hartnn):

nopes

OpenStudy (anonymous):

well, its like distance formula except it has r squared in it

OpenStudy (anonymous):

and no sq. root

OpenStudy (anonymous):

64+64=128=r sq.

hartnn (hartnn):

put (h,k)=(6,8) in that

OpenStudy (anonymous):

sq. root 128

OpenStudy (anonymous):

ahh

OpenStudy (anonymous):

whose center is at 6,8

OpenStudy (anonymous):

I thought the center was at the origin

OpenStudy (anonymous):

64+64=r sq.

OpenStudy (anonymous):

r= sq. root 128

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

what next?

hartnn (hartnn):

nopes, u did not put h,k = 6,8 in this. (x-h)^2 + (y-k)^2 = r^2

OpenStudy (anonymous):

yeah I did

OpenStudy (anonymous):

(0-8) quantity squared=64

OpenStudy (anonymous):

ahh

OpenStudy (anonymous):

64+36=r sq.

OpenStudy (anonymous):

r=10

OpenStudy (anonymous):

l

OpenStudy (anonymous):

k

OpenStudy (anonymous):

so the radius equals 10

OpenStudy (anonymous):

(10,0)?

OpenStudy (anonymous):

or just 10?

OpenStudy (anonymous):

so (16,8) is a possibility

OpenStudy (anonymous):

so is (4,-8)

hartnn (hartnn):

radius is 10, but i would say u still did NOT put h,k = 6,8 in (x-h)^2 + (y-k)^2 = r^2

OpenStudy (anonymous):

as well as (6,8)

OpenStudy (anonymous):

I mean

OpenStudy (anonymous):

(6,-2)

OpenStudy (anonymous):

so we have two answers left, (12,0) and (-2,12)

OpenStudy (anonymous):

where did I go wrong?

OpenStudy (anonymous):

36+64=100

OpenStudy (anonymous):

r=10

OpenStudy (anonymous):

I plugged the numbers in the formula

hartnn (hartnn):

can i request u something ? please put h,k = 6,8 in (x-h)^2 + (y-k)^2 = r^2

OpenStudy (anonymous):

h=36

OpenStudy (anonymous):

after quantity squared

OpenStudy (anonymous):

k=64 quantity squared

OpenStudy (anonymous):

wait a minute

OpenStudy (anonymous):

ahh

OpenStudy (anonymous):

(-2-8) sq. +(12-6) sq=r sq.

OpenStudy (anonymous):

D. is incorrect

OpenStudy (anonymous):

is that right?

hartnn (hartnn):

let me put h,k = 6,8 in and r=10 (x-h)^2 + (y-k)^2 = r^2 put h,k = 6,8 in (x-6)^2 + (y-8)^2 = 10^2 did u get this ?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

r=10

OpenStudy (anonymous):

D. is incorrect because r does not equal 10

OpenStudy (anonymous):

so D. is the answer for the problem

OpenStudy (anonymous):

ty for teaching me that formula

OpenStudy (anonymous):

It will be useful for me in the future

hartnn (hartnn):

yes, it will be welcome :)